OpenStudy (anonymous):

grandpa's age is 6 yrs less than 6 times junior's age. the sum of their age is 78. find each of their ages

OpenStudy (anonymous):

ok, grandma's represents y, jr is x. you have to equations, y = 6x-6 (grandma's age) and x + y = 78. here, you have to do substitution. so... y = 78 - x. from there you set that equal to the other equation, so 78 - x = 6x - 6. you then solve for x (junior's age) and plug that in in order to find grandma's age. hope that helps!